2.given TC=400-50Q+Q2 P=80 .find the optimal quantity produced and sold by the firm so that they are maximizing profits


1
Expert's answer
2021-12-14T09:54:54-0500

Solution:

Profit maximizing quantity is where MR = MC

MR = P = 80

MC = "\frac{\partial TR} {\partial Q}" = 50 + 2Q

Set MR = MC

80 = 50 + 2Q

80 – 50 = 2Q

30 = 2Q

Q = 15

The optimal quantity produced and sold by the firm so that they are maximizing = 15 units